# Population dynamics simulations of large deviations for three subclasses of the Kardar-Parisi-Zhang universality class

### v1

@@ II.1 Model
The studied model, TASEP, is the strong asymmetry limit of the asymmetric simple exclusion process (ASEP) [12, 20], a paradigmatic model of stochastic particle transport on a one-dimensional lattice. In TASEP, particles hop stochastically to their right neighbor sites at a constant rate (set to be $1$) if the target sites are empty (Fig. 1(b)). The only interaction between particles is volume exclusion, which prevents them from hopping to sites that are already occupied; otherwise particles undergo stochastic hopping independently. This model is known (and mathematically proven) to be in the KPZ class. Indeed, by replacing occupied and unoccupied sites with downward and upward slopes, respectively, one can map particle configurations to interface profiles (Fig. 1(c)-(e)) and each hopping of a particle corresponds to a local growth of the interface at the corresponding position. The increment of the interface height $h(x,t)$ is given by the time-integrated current of particles, i.e., the total number of particles that have passed position $x$ rightward until time $t$. In the following, we only analyze $h$ at $x=0$, which is the center of the lattice of size $N$ (see Fig. 1(c)-(e) again). With this, we define the rescaled height fluctuation $\delta h$ by
@@
In the present work, we study the three representative subclasses of the one-dimensional KPZ class, corresponding to the step, flat, and stationary initial conditions (Fig. 1(c)-(e)). The step initial condition is a configuration where the left half of the system is fully occupied and the right half is fully empty (Fig. 1(c)), and typical fluctuations in this case are known to show the GUE Tracy-Widom distribution [25, 8, 37]. The flat initial condition refers to a configuration where occupied and unoccupied sites are alternating (Fig. 1(d)) and typical fluctuations show the GOE Tracy-Widom distribution. For the stationary initial condition, an initial configuration is drawn from the Bernouilli distribution with density $\rho=1/2$ (Fig. 1(e)), i.e., each site is set to be occupied or unoccupied with equal probability. Here we also impose the constraint that the total number of occupied sites is half of the number of sites, $N$. Practically, we choose $N/2$ out of $N$ sites randomly and put particles therein. Typical fluctuations in this stationary subclass are known to be characterized by the Baik-Rains distribution.
@@ V Stationary initial condition
Figure 5: Simulation results for the stationary initial condition with $N=10000$ and $N_{\mathrm{cl}}=10000$ from 10001000 realizations (unless otherwise stipulated). (a)(b) The absolute value of CGF $|\lambda(k,t)|$ (a) and its local exponent $\frac{\mathrm{d}\ln|\lambda(k,t)|}{\mathrm{d}\ln t}$ (b) for negative bias $k<0$. In (a), simulation data for the stationary initial condition (symbols) are compared with the exact solution for the step initial condition (solid lines), Eq. (10), or the prediction for typical fluctuations for the stationary initial condition (dashed line), Eq. (21). (c) The scaled CGF $\mu_{-}(k)$ for the negative tail. Simulation data of $\lambda(k,t)/t$ for the stationary initial condition (symbols) are compared with the exact solution for the step initial condition (line), Eq. (10). The values of $t$ used are $1000,1000,575,184,75$ from the leftmost to the rightmost datapoints. Inset: the ratio of $\lambda(k,t)/t$ for the stationary initial condition to the exact solution of $\mu^{\mathrm{step}}_{-}(k)$ for the step initial condition, with k=0.464k=-0.464 (blue), 1-1 (green), and 4.64-4.64 (red). (d) An example of time evolution of interface profiles for k=1k=-1, here with $N=1024$. (e)(f) The CGF $\lambda(k,t)$ (e) and its local exponent $\frac{\mathrm{d}\ln\lambda(k,t)}{\mathrm{d}\ln t}$ (f) for positive bias $k>0$. In (e), simulation data for the stationary initial condition (symbols) are compared with the exact solution for the step initial condition (solid lines), Eq. (11), or the prediction for typical fluctuations for the stationary initial condition (dashed line), Eq. (16). Inset: the scaled CGF $\mu_{+}(k/t)$ for the positive tail. Simulation data of $\lambda(k,t)/t^{2}$ for the stationary initial condition (symbols) are plotted against $k/t$ and compared with the exact solution for the step initial condition (line), Eq. (11). Inset: the ratio of $\lambda(k,t)/t^{2}$ for the stationary initial condition to the exact solution of $\mu^{\mathrm{step}}_{+}(k/t)$ for the step initial condition. (h) An example of time evolution of interface profiles for $k=1$, here with $N=1024$.
@@
Figure 5 shows the numerical results of our population dynamics simulations for the stationary case. Since the stationary initial condition is stochastic, here we repeated simulations from different initial realizations $1000$ times and took average of $\lambda(k,t)$ over all realizations. Then the CGF $\lambda(k,t)$ shows crossover from the typical fluctuation regime [Eq. (21)], $\lambda(k,t)\sim t^{2/3}$ which is characteristic of the stationary case, to the large deviation regimes [Eq. (24)] [Fig. 5(a)(b)(e)(f)]. Due to the larger exponent value of the typical fluctuation regime, we were able to reach the large deviation regimes only for relatively large values of $|k|$, compared to the step and flat cases. This prevented us from testing Eq. (22) for $k\to-0$. For relatively large $|k|$ for which we could reach the large deviation regime, our data indicate $\mu^{\mathrm{stat}}_{-}(k)\simeq\mu^{\mathrm{step}}_{-}(k)$ [Fig. 5(c)], that is, the scaled CGF $\mu_{-}(k)$ remains identical to that for the step initial condition. Similarly to the flat case, this coincidence can be interpreted as a result of the spontaneous formation of wedge interface profiles for $k<0$ [Fig. 4(d)]. Regarding the positive tail, we confirm $\mu_{+}(k/t)\sim(k/t)^{3/2}$ [Fig. 5(g)], but we were unable to demonstrate the data collapse with different $k$ (inset), most likely due to the limited time range $t\leq 10^{3}$ we had to work with for repeating many realizations. It is worthwhile to accomplish this computational task in future, in view of the conjecture for the KPZ equation [7, 2] that $\Phi_{\pm}^{(\mp)}(z)$, and equivalently $\mu_{\pm}^{(\mp)}(k)$ (taking the signs in the same order), are identical between the step and stationary initial conditions.

### v2

@@ II.1 Model
The studied model, TASEP, is the strong asymmetry limit of the asymmetric simple exclusion process (ASEP) [12, 20], a paradigmatic model of stochastic particle transport on a one-dimensional lattice. In TASEP, particles hop stochastically to their right neighbor sites at a constant rate (set to be 11) if the target sites are empty (Fig. 1(b)). The only interaction between particles is volume exclusion, which prevents them from hopping to sites that are already occupied; otherwise particles undergo stochastic hopping independently. This model is known (and mathematically proven) to be in the KPZ class. Indeed, by replacing occupied and unoccupied sites with downward and upward slopes, respectively, one can map particle configurations to interface profiles (Fig. 1(c)-(e)) and each hopping of a particle corresponds to a local growth of the interface at the corresponding position. The increment of the interface height $h(x,t)$ is given by the time-integrated current of particles, i.e., the total number of particles that have passed position $x$ rightward until time $t$. In the following, we only analyze $h$ at $x=0$, which is the center of the lattice of size NN (see Fig. 1(c)-(e) again). With this, we define the rescaled height fluctuation $\delta h$ by
@@
In the present work, we study the three representative subclasses of the one-dimensional KPZ class, corresponding to the step, flat, and stationary initial conditions (Fig. 1(c)-(e)). The step initial condition is a configuration where the left half of the system is fully occupied and the right half is fully empty (Fig. 1(c)), and typical fluctuations in this case are known to show the GUE Tracy-Widom distribution [25, 8, 37]. The flat initial condition refers to a configuration where occupied and unoccupied sites are alternating (Fig. 1(d)) and typical fluctuations show the GOE Tracy-Widom distribution. For the stationary initial condition, an initial configuration is drawn from the Bernouilli distribution with density $\rho=1/2$ (Fig. 1(e)), i.e., each site is set to be occupied or unoccupied with equal probability. Here we also impose the constraint that the total number of occupied sites is half of the number of sites, NN. Practically, we choose $N/2$ out of NN sites randomly and put particles therein. Typical fluctuations in this stationary subclass are known to be characterized by the Baik-Rains distribution.
@@ V Stationary initial condition
Figure 5: Simulation results for the stationary initial condition with $N=10000$ and $N_{\mathrm{cl}}=1000$ from 10001000 realizations (unless otherwise stipulated). (a)(b) The absolute value of CGF $|\lambda(k,t)|$ (a) and its local exponent $\frac{\mathrm{d}\ln|\lambda(k,t)|}{\mathrm{d}\ln t}$ (b) for negative bias $k<0$. In (a), simulation data for the stationary initial condition (symbols) are compared with the exact solution for the step initial condition (solid lines), Eq. (10), or the prediction for typical fluctuations for the stationary initial condition (dashed line), Eq. (21). (c) The scaled CGF $\mu_{-}(k)$ for the negative tail. Simulation data of $\lambda(k,t)/t$ for the stationary initial condition (symbols) are compared with the exact solution for the step initial condition (line), Eq. (10). The values of $t$ used are $1000,1000,575,184,75$ from the leftmost to the rightmost datapoints. Inset: the ratio of $\lambda(k,t)/t$ for the stationary initial condition to the exact solution of $\mu^{\mathrm{step}}_{-}(k)$ for the step initial condition, with k=0.464k=-0.464 (blue), 1-1 (green), and 4.64-4.64 (red). (d) An example of time evolution of interface profiles for k=1k=-1, here with $N=1024$ and $N_{\mathrm{cl}}=10000$. (e)(f) The CGF $\lambda(k,t)$ (e) and its local exponent $\frac{\mathrm{d}\ln\lambda(k,t)}{\mathrm{d}\ln t}$ (f) for positive bias $k>0$. In (e), simulation data for the stationary initial condition (symbols) are compared with the exact solution for the step initial condition (solid lines), Eq. (11), or the prediction for typical fluctuations for the stationary initial condition (dashed line), Eq. (16). Inset: the scaled CGF $\mu_{+}(k/t)$ for the positive tail. Simulation data of $\lambda(k,t)/t^{2}$ for the stationary initial condition (symbols) are plotted against $k/t$ and compared with the exact solution for the step initial condition (line), Eq. (11). Inset: the ratio of $\lambda(k,t)/t^{2}$ for the stationary initial condition to the exact solution of $\mu^{\mathrm{step}}_{+}(k/t)$ for the step initial condition. (h) An example of time evolution of interface profiles for $k=1$, here with $N=1024$ and $N_{\mathrm{cl}}=10000$.
@@
Figure 5 shows the numerical results of our population dynamics simulations for the stationary case. Since the stationary initial condition is stochastic, here we repeated simulations from different initial realizations 10001000 times and took average of $\lambda(k,t)$ over all realizations. Then the CGF $\lambda(k,t)$ shows crossover from the typical fluctuation regime [Eq. (21)], $\lambda(k,t)\sim t^{2/3}$ which is characteristic of the stationary case, to the large deviation regimes [Eq. (24)] [Fig. 5(a)(b)(e)(f)]. Due to the larger exponent value of the typical fluctuation regime, we were able to reach the large deviation regimes only for relatively large values of $|k|$, compared to the step and flat cases. This prevented us from testing Eq. (22) for $k\to-0$. For relatively large $|k|$ for which we could reach the large deviation regime, our data indicate $\mu^{\mathrm{stat}}_{-}(k)\simeq\mu^{\mathrm{step}}_{-}(k)$ [Fig. 5(c)], that is, the scaled CGF $\mu_{-}(k)$ remains identical to that for the step initial condition. Similarly to the flat case, this coincidence can be interpreted as a result of the spontaneous formation of wedge interface profiles for $k<0$ [Fig. 4(d)]. Regarding the positive tail, we confirm $\mu_{+}(k/t)\sim(k/t)^{3/2}$ [Fig. 5(g)], but we were unable to demonstrate the data collapse with different $k$ (inset), most likely due to the limited time range $t\leq 10^{3}$ we had to work with for repeating many realizations. It is worthwhile to accomplish this computational task in future, in view of the conjecture for the KPZ equation [7, 2] that $\Phi_{\pm}^{(\mp)}(z)$, and equivalently $\mu_{\pm}^{(\mp)}(k)$ (taking the signs in the same order), are identical between the step and stationary initial conditions.